# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ Abstract.
Relevant physical phenomena are described by nonlinear Schrödinger equations with non-vanishing conditions at infinity. This paper investigates the respective 2D and 3D Cauchy problems. Local well-posedness in the energy space for energy-subcritical nonlinearities, merely satisfying Kato-type assumptions, is proven, providing the analogue of the well-established local $H^{1}$-theory for solutions vanishing at infinity. The critical nonlinearity will be simply a byproduct of our analysis and the existing literature. Under an assumption that prevents the onset of a Benjamin-Feir type instability, global well-posedness in the energy space is proven for a) non-negative Hamiltonians, b) sign-indefinite Hamiltonians under additional assumptions on the zeros of the nonlinearity, c) generic nonlinearities and small initial data. The cases b) and c) only concern 3D.
@@ 1. Introduction
(1.3) $\mathcal{H}(\psi)=\int_{{\mathbf{R}}^{d}}\frac{1}{2}|\nabla\psi|^{2}+F(|\psi|^{2})\mathrm{d}x,\quad\text{with}\quad F(\rho)=\int_{\rho_{0}}^{\rho}f(r)\mathrm{d}r.$
@@
In the first part of the paper, we establish local well-posedness in the energy space for (1.1)-(1.2) with energy-subcritical nonlinear potentials $f$ under Kato-type [39] regularity assumptions. The continuity of the solution map is proven with respect to the topology of the (curved) energy space and not only in affine spaces.
Second, global well-posedness is proven, provided that $f^{\prime}(1)>0$, see Assumption 1.5 below. Specifically, global well-posedness is shown for sign-definite total energies and $d=2,3$, and for sign-indefinite total energies and $d=3$ under suitable additional assumptions on $f$ and the decay of the initial data at infinity or in alternative for small initial data.
@@
For general defocusing nonlinearities, including the nonlinearities considered in Assumption 1.5 below, the existence of sub-sonic traveling waves is investigated in [54, 17]. Non-existence in the super-sonic regime is shown in [53]. For $d=2$, traveling waves exist for any, and in particular arbitrarily small energy ruling out scattering, while for $d=3$ there is an energy threshold below of which no traveling waves exist. We remark that the assumptions given in the paper of [54, 17] are strongly related with our assumptions on the nonlinear potential $f$. The stability of multi-dimensional traveling waves is addressed in [16, 52], stationary bubbles and their stability in [19]. Transverse instability is studied in [50]. The GP equation admits vortex solutions with infinite energy, see [59, 11] and [65, 30] as well as references therein for stability properties.
@@ 1.1. Previous well-posedness results
Local existence of solutions to GP in Zhidkov spaces has been investigated in [66, 68] for $d=1$ and [21] for the multi-dimensional case. For $d=1$, global well-posedness of GP in the energy space is shown in [66] and has recently been proven for fractional Sobolev and low regularity in [48, 49]. While the energy space for GP for $d=1$ coincides with the set of functions in the Zhidkov space, such that $|\psi|^{2}-1\in L^{2}({\mathbf{R}})$, this identification does not hold true in the multi-dimensional case, see [23] and Section 2 below. The GP is well-posed in $1+H^{1}({\mathbf{R}}^{d})$ for $d=2,3$ [10]. Global well-posedness in $1+H^{s}({\mathbf{R}}^{3})$ with $s\in(5/6,1)$ is proven in [57]. However, the space $1+H^{1}({\mathbf{R}}^{d})$ is strictly smaller than the natural energy space $\mathbb{E}({\mathbf{R}}^{d})$, see (1.9) below. In fact, there exist traveling waves for GP in the energy space that do not belong to $1+L^{2}({\mathbf{R}}^{d})$, see [29]. Global well-posedness in the energy space for the multi-dimensional GP has been introduced in the seminal paper [23]. One of the major novelties of [23] consists in the precise characterization of the energy space as complete metric space and the action of the free propagator on the energy space. A more general class of defocusing and energy-subcritical $C^{3}$-nonlinearities has been considered in [22] with subsequent improvement to $C^{2}$-nonlinearities [56]. In [22, 56], the respective authors crucially rely on a smooth decomposition of wave-functions in the energy space. Global well-posedness is proven in affine spaces determined by this decomposition which requires the aforementioned regularity assumptions and precise growth conditions for $f$. The result in the affine spaces then implies existence and uniqueness in the energy space. The cubic-quintic equation being energy-critical is studied in [44, 42, 43].
In [12], global existence of unique mild solutions to (1.1) with a logarithmic nonlinearity is introduced.
@@ 1.2. Local well-posedness results
    $f\in C([0,\infty))\cap C^{1}((0,\infty))$ such that $f(1)=0$,
@@
(1.4) $f(|\psi|^{2})=\lambda(|\psi|^{2\alpha}-1),\qquad\text{with}\quad\lambda=\pm 1\,\,\text{and}\,\,\begin{cases}\alpha>0\quad&d=2,\\ 0<\alpha<2\quad&d=3.\end{cases}$
@@
(1.8) $\mathbb{E}_{\mathrm{GL}}$ =$=\{\psi\in L_{\operatorname{loc}}^{1}({\mathbf{R}}^{d}):\mathcal{H}(\psi)<+\infty\}$
=$=\{\psi\in L_{\operatorname{loc}}^{1}({\mathbf{R}}^{d}):\nabla\psi\in L^{2}({\mathbf{R}}^{d}),|\psi|^{2}-1\in L^{2}({\mathbf{R}}^{d})\}.$
@@
It is straightforward to see that $\mathbb{E}\subset\mathbb{E}_{\mathrm{GL}}$. However, as it will be clear later, see Lemmata 2.6 and 2.8, the two spaces $\mathbb{E}$ and $\mathbb{E}_{\mathrm{GL}}$ turn out to be equivalent. Working in $\mathbb{E}$ rather than $\mathbb{E}_{\mathrm{GL}}$ is more convenient in several aspects when dealing with a general class of nonlinearities $f$ satisfying Assumption 1.1.
@@
(1.11) $X^{1}({\mathbf{R}}^{d})=\{\psi\in L^{\infty}({\mathbf{R}}^{d}):\nabla\psi\in L^{2}({\mathbf{R}}^{d})\}.$
@@
$\|\psi\|_{X^{1}+H^{1}}=\inf\left\{\|\psi_{1}\|_{X_{1}}+\|\psi_{2}\|_{H^{1}}\,:\,\psi=\psi_{1}+\psi_{2}\right\}.$
@@
Let $d=2,3$. Let $f$ be such that Assumption 1.1 is satisfied, then (1.1) is locally well-posed in the energy space $\mathbb{E}({\mathbf{R}}^{d})$. More precisely,
@@
    for any $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$, there exists a maximal time of existence $T^{\ast}>0$ and a unique solution $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{d}))$ with initial data $\psi(0)=\psi_{0}$. The following blow-up alternative holds namely, either $T_{\ast}=\infty$ or
@@
    the following identity holds $\mathcal{H}(\psi)(t)=\mathcal{H}(\psi_{0}),$ for all $t\in[0,T^{\ast})$;
@@
Note that (2) of Theorem 1.3 states that $\psi$ and $\psi_{0}$ share the same far-field behavior, i.e. belong to the same connected component of $\mathbb{E}({\mathbf{R}}^{d})$ for all $t\in[0,T^{\ast})$, see Remark 2.3 and 2.4. Moreover, it can be shown that the nonlinear flow $\psi-\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}$ belongs to the full range of Strichartz spaces, see Proposition 3.2 and 4.1 for $d=2,3$ respectively. The precise notion of continuous dependence on the initial data is given in Proposition 3.2 and 4.1. The topological structure of the metric space $(\mathbb{E}({\mathbf{R}}^{d}),d_{\mathbb{E}})$ differs for $d=2$ and $d=3$, see [23, 24]. For $d=3$, the energy space $\mathbb{E}({\mathbf{R}}^{3})$ has an affine structure; if $\psi\in\mathbb{E}({\mathbf{R}}^{3})$ then $\psi=c+v$ for some $c\in S^{1}$, $v\in\dot{H}^{1}({\mathbf{R}}^{3})$. For $d=2$, unbounded phase oscillations may occur at spatial infinity that rule out to characterize the connected components by a constant $c\in S^{1}$. The space $(\mathbb{E}({\mathbf{R}}^{2}),d_{\mathbb{E}})$ is not separable. Given its relevance for the well-posedness theory, this question is going to be addressed in detail in Section 2. In particular, one may introduce a weaker topology that restores separability and connectedness. Note that this affine structure of the energy space is available for higher dimensions $d\geq 4$ to which our approach adapts. As $\mathbb{E}({\mathbf{R}})\subset X^{1}({\mathbf{R}})$, the local well-posedness theory simplifies for $d=1$. We expect our approach to extend to $d=1$. Previous results [21, 22, 25] do not cover the full generality of Assumption 1.1.
@@
Namely, for any $r,R>0$ and $\psi_{0}^{\ast}\in\mathbb{E}({\mathbf{R}}^{d})$ such that $\mathcal{E}(\psi_{0}^{\ast})\leq R$ let $\mathcal{O}_{r}:=\{\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d}):d(\psi_{0},\psi_{0}^{\ast})\leq r\}$. Then, there exists $T^{\ast}(\mathcal{O}_{r})>0$ such that $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{d}))$ for all initial data $\psi(0)=\psi_{0}\in\mathcal{O}_{r}$. Moreover, for any $0<T<T^{\ast}(\mathcal{O}_{r})$ there exists $C>0$ such that for any $\psi_{1},\psi_{2}\in C([0,T];\mathbb{E}({\mathbf{R}}^{d}))$ with initial data $\psi_{0}^{1},\psi_{0}^{2}\in\mathcal{O}_{r}$ it holds
@@ 1.3. Global well-posedness results
This assumption yields that $F$ achieves a local minimum for the constant solution $|\psi|^{2}=1$. In nonlinear optics, this requirement is made in physical literature in order to prevent the onset of modulational instability, also referred to as Benjamin-Feir instability [6], of the constant equilibrium solution, i.e. the continuous wave background [45, 58].
@@
with $a>0$, $20<\alpha_{1}<$ 2 and where $g$ satisfies Assumptions 1.1 (K1) and (K2) for some $0\leq\alpha_{2}<\alpha_{1}$. In addition, $F$ is such that $F(\rho)>0$ for all $\rho>1$.
@@
The assumption on the roots of $F$ allows for physically relevant nonlinearities to be studied. It appears from the physics literature [5, 46, 58] that in relevant applications the largest root of $F$ corresponds to the far-field behavior $\rho_{0}=1$ and constitutes a local minimum of $F$ which is linked to preventing modulational instability of the continuous background wave [45, 58]. To obtain global existence, we rely on the aforementioned affine structure of the energy space $\mathbb{E}({\mathbf{R}}^{3})$ and require that $\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{3})$ while $v_{0}\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ only yields $|v_{0}|^{2}+2\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{2})$. An exponential bound on $\operatorname{Re}(\overline{c}v)(t)\in L^{2}({\mathbf{R}}^{3})$ is derived which compensates for the lack of the conserved mass due to the non-trivial farfield. The result of Theorem 1.7 remains valid if the assumption $\operatorname{Re}(\overline{c}v_{0})\in L^{2}$ is replaced by a smallness assumption on $\mathcal{H}(\psi_{0})$ and $\|\nabla\operatorname{Re}(\overline{c}v_{0})\|_{L^{2}}^{2}$ depending only on the second largest positive root of $F$, see Remark 4.12.
Finally, we consider the general scenario for $F$ being such that Assumption 1.5 is satisfied but $F$ possibly unbounded from below, e.g. in the case of competing power-law nonlinearities with the focusing one dominant at large intensities.
@@
To the best of the authors’ knowledge, global results for (1.1) in $d=2,3$ are in general not available in the literature in the case of non sign-definite total energies unless the nonlinear potential energy density is assumed to be bounded from below and in addition more regularity on $f$ [22], or in the cubic-quintic case a condition on $\operatorname{Re}(v)$ like the ones mentioned, cf. [44, 42] are assumed. In [43] small-data global well-posedness for cubic-quintic nonlinearities is proven also in the case where the quintic nonlinearity is focusing. In [55] the authors consider for $d=1,2$ nonlinear potentials energy density unbounded from below for specific regular energy subcritical nonlinearities and prove small data global existence for solutions of the form $\psi=1+u$ in tailored function spaces.
The main steps of our approach are briefly sketched. First, we identify the suitable mathematical setting for our analysis, namely the energy space $\mathbb{E}$, see (1.9). We crucially rely on the fact that $(\mathbb{E},d_{\mathbb{E}})$ is a complete metric space as well as the properties of the free propagator introduced in [23, 24]. The Hamiltonian $\mathcal{H}$ is well-defined for functions in $\mathbb{E}$. While wave-functions in $d=3$ can be decomposed as $\psi=c+v$ with $|c|=1,c\in{\mathbf{C}}$ and $v\in\dot{H}^{1}({\mathbf{R}}^{3})$, for $d=2$ the wave-functions may exhibit unbounded oscillations of the phase at spatial infinity. This motivates to treat separately the well-posedness problem for $d=2,3$. In both cases, we show local existence of a solution in the affine space $\psi=\psi_{0}+H^{1}({\mathbf{R}}^{d})$ by a perturbative Kato-type argument [39] and also [14, Chapter 4]. Subsequently, uniqueness in $C([0,T];\mathbb{E}({\mathbf{R}}^{d}))$ is proven. The fixed-point argument only provides continuous dependence with respect to perturbations in the space $\psi_{0}+H^{1}({\mathbf{R}}^{d})$. The proof of continuous dependence on the initial data with respect to the topology induced by the metric $d_{\mathbb{E}}$ requires additional estimates and differs in a substantial way from the $H^{1}$-well-posedness theory for NLS-equations with vanishing conditions at infinity. This is due to the non-integrability of wave-functions and the intricate topological structure of the energy space linked to the far-field behavior including oscillations of the phase and the low regularity of the nonlinearity. Global well-posedness is shown relying on the conservation of the Hamiltonian $\mathcal{H}$.
@@
For the 3D energy-critical critical quintic equation, one may proceed as described in Section 1.4.
@@
Beyond the mentioned power-type nonlinearities, the following are examples of physically relevant nonlinearties and far-field (1.2):
@@
    saturated nonlinearities $f(\rho)=\frac{\rho}{1+\gamma\rho}-\frac{1}{1+\gamma}$ with $\gamma>0$, see for instance [61, Chapter 9.3] and references therein,
@@
    exponential nonlinearities $f(\rho)=(\mathrm{e}^{-\gamma}-\mathrm{e}^{-\gamma\rho})$ with $\gamma>0$ [61, Chapter 9.3],
@@
    transiting nonlinearities of the form $f(\rho)=2\rho\left(1+\alpha\tanh\left(\gamma(\rho^{2}-1\right)\right)$ occurring in nonlinear optics [58, Section VI],
@@
The cubic-quintic equation (1.18) falls within (1) of the aforementioned list and is also recovered in the small amplitude approximation of (2) and (3) of the above examples [61, Chapter 9.3].

### v4

@@ Abstract.
Relevant physical phenomena are described by nonlinear Schrödinger equations with non-vanishing conditions at infinity. This paper investigates the respective 2D and 3D Cauchy problems. Local well-posedness in the (curved) energy space for energy-subcritical nonlinearities, merely satisfying Kato-type assumptions, is proven, providing the analogue of the well-established local $H^{1}$-theory for solutions vanishing at infinity. The critical nonlinearity will be simply a byproduct of our analysis and the existing literature. Under an assumption that prevents the onset of a Benjamin-Feir type instability, global well-posedness in the energy space is proven for a) non-negative Hamiltonians, b) sign-indefinite Hamiltonians under additional assumptions on the zeros of the nonlinearity, c) generic nonlinearities and small initial data. The cases b) and c) only concern the 3D case.
@@ 1. Introduction
(1.3) $\mathcal{H}(\psi)=\int_{{\mathbf{R}}^{d}}\frac{1}{2}|\nabla\psi|^{2}+F(|\psi|^{2})\mathrm{d}x,\quad\text{with}\quad F(\rho)=\int_{1}^{\rho}f(r)\mathrm{d}r.$
@@
In the first part of the paper, we establish local well-posedness in the energy space for (1.1)-(1.2) with energy-subcritical nonlinear potentials $f$ under Kato-type [42] regularity assumptions. The continuity of the solution map is proven with respect to the topology of the (curved) energy space and not only in affine spaces. The choice of a suitable functional framework plays a crucial role for the stability analysis of particular solutions [17, 33]. Indeed, as pointed out in [27] and [20, Remark 1.2] the constant solution with $|c|=1$ is linearly unstable in the affine space $1+H^{1}({\mathbf{R}}^{d})$ while orbitally stable in the energy space for $d=1,2$. A similar result holds for the Ginzburg-Landau vortex of degree one [33].
Second, global well-posedness is proven, provided that $f^{\prime}(1)>0$, see Assumption 1.5 below. Specifically, global well-posedness is shown for sign-definite total energies and $d=2,3$, and for sign-indefinite total energies and $d=3$ under suitable additional assumptions on $f$ and the decay of the initial data at infinity or alternatively for small initial data.
@@
For general defocusing nonlinearities, including the nonlinearities considered in Assumption 1.5 below, the existence of sub-sonic traveling waves is investigated in [58, 17]. Non-existence in the super-sonic regime is shown in [57]. For $d=2$, traveling waves exist for any, and in particular arbitrarily small energy ruling out scattering, while for $d=3$ there is an energy threshold below which no traveling waves exist. We remark that the assumptions given in [58, 17] are strongly related with our assumptions on the nonlinear potential $f$. The stability of multi-dimensional traveling waves is addressed in [16, 56] and stationary bubbles and their stability in [21]. Transverse instability is studied in [54]. The GP equation admits vortex solutions with infinite energy, see [63, 11] and [69, 33] as well as references therein for stability properties.
@@ 1.1. Previous well-posedness results
Local existence of solutions to GP in Zhidkov spaces has been investigated in [71, 73] for $d=1$ and [24] for the multi-dimensional case. For $d=1$, GP is known to be completely integrable [70]. The global well-posedness of GP in the energy space is shown in [71] and has recently been proven for fractional Sobolev and low regularity in [51, 52]. While the energy space for GP for $d=1$ coincides with the set of functions in the Zhidkov space, such that $|\psi|^{2}-1\in L^{2}({\mathbf{R}})$, this identification does not hold true in the multi-dimensional case, see [26] and Section 2 below. The GP is well-posed in $1+H^{1}({\mathbf{R}}^{d})$ for $d=2,3$ [10]. Global well-posedness in $1+H^{s}({\mathbf{R}}^{3})$ with $s\in(5/6,1)$ is proven in [61]. However, the space $1+H^{1}({\mathbf{R}}^{d})$ is strictly smaller than the natural energy space $\mathbb{E}({\mathbf{R}}^{d})$, see (1.9) below. In fact, there exist traveling waves for GP in the energy space that do not belong to $1+L^{2}({\mathbf{R}}^{d})$, see [32]. Global well-posedness in the energy space for the multi-dimensional GP has been introduced in the seminal paper [26]. One of the major novelties of [26] consists in the precise characterization of the energy space as complete metric space and the action of the free propagator on the energy space. A more general class of defocusing and energy-subcritical $C^{3}$-nonlinearities has been considered in [25] with subsequent improvement to $C^{2}$-nonlinearities [60]. In [25, 60], the respective authors crucially rely on a smooth decomposition of wave-functions in the energy space. Global well-posedness is proven in affine spaces determined by this decomposition which requires the aforementioned regularity assumptions and precise growth conditions for $f$. The result in the affine spaces then implies existence and uniqueness in the energy space. The cubic-quintic equation being energy-critical is studied in [47, 45, 46].
In [12], global existence of unique mild solutions to (1.1)-(1.2) with a logarithmic nonlinearity is introduced.
@@ 1.2. Local well-posedness results
    $f\in C([0,\infty))\cap C^{1}((0,\infty))$ and is such that $f(1)=0$,
@@
(1.4) $f(|\psi|^{2})=\lambda(|\psi|^{2\alpha}-1),\qquad\text{with}\quad\lambda=\pm 1\,\,\text{and}\,\,\begin{cases}\alpha>0\quad&\text{for}\,d=2,\\ 0<\alpha<2\quad&\text{for}\,d=3.\end{cases}$
@@
(1.8) $\mathbb{E}_{\mathrm{GL}}$ $=\{\psi\in L_{\operatorname{loc}}^{1}({\mathbf{R}}^{d}):\mathcal{H}(\psi)<+\infty\}$
$=\{\psi\in L_{\operatorname{loc}}^{1}({\mathbf{R}}^{d}):\nabla\psi\in L^{2}({\mathbf{R}}^{d}),|\psi|^{2}-1\in L^{2}({\mathbf{R}}^{d})\}.$
@@
As $||\psi|-1|\leq||\psi|^{2}-1|$ it follows that $\mathbb{E}_{\mathrm{GL}}\subset\mathbb{E}$ and the converse inclusion is straightforward to check, see Lemma 2.5. Working in $\mathbb{E}$ rather than $\mathbb{E}_{\mathrm{GL}}$ is more convenient in several aspects when dealing with a general class of nonlinearities $f$ satisfying Assumption 1.1.
@@
(1.11) $X^{1}({\mathbf{R}}^{d})=\{\psi\in L^{\infty}({\mathbf{R}}^{d}):\nabla\psi\in L^{2}({\mathbf{R}}^{d})\},\quad\|\psi\|_{X^{1}({\mathbf{R}}^{d})}:=\|\psi\|_{L^{\infty}({\mathbf{R}}^{d})}+\|\nabla\psi\|_{L^{2}({\mathbf{R}}^{d})}.$
@@
$\|\psi\|_{X^{1}+H^{1}}=\inf\left\{\|\phi\|_{X_{1}}+\|u\|_{H^{1}}\,:\,\psi=\phi+u\right\}.$
Note, that the two metric spaces $\mathbb{E}$ and $\mathbb{E}_{\mathrm{GL}}$ turn out to be equivalent, see Lemmata 2.6 and 2.8 below.
@@
Let $d=2,3$ and let $f$ be as in Assumption 1.1. Then (1.1) is locally well-posed in the energy space $\mathbb{E}({\mathbf{R}}^{d})$. More precisely,
@@
    for any $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$, there exist a maximal time of existence $T^{\ast}>0$ and a unique solution $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{d}))$ with initial data $\psi(0)=\psi_{0}$. The following blow-up alternative holds namely, either $T_{\ast}=\infty$ or
@@
    the identity $\mathcal{H}(\psi)(t)=\mathcal{H}(\psi_{0})$ holds for all $t\in[0,T^{\ast})$;
@@
Note that (2) of Theorem 1.3 states that $\psi$ and $\psi_{0}$ share the same far-field behavior, i.e. they belong to the same connected component of $\mathbb{E}({\mathbf{R}}^{d})$ for all $t\in[0,T^{\ast})$, see Remarks 2.3 and 2.4. Moreover, it can be shown that the nonlinear flow $\psi-\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}$ belongs to the full range of Strichartz spaces, see Proposition 3.2 and 4.1 for $d=2,3$ respectively. The precise notion of continuous dependence on the initial data is given in Proposition 3.2 and 4.1. The topological structure of the metric space $(\mathbb{E}({\mathbf{R}}^{d}),d_{\mathbb{E}})$ differs for $d=2$ and $d=3$, see [26, 27]. For $d=3$, the energy space $\mathbb{E}({\mathbf{R}}^{3})$ has an affine structure; if $\psi\in\mathbb{E}({\mathbf{R}}^{3})$ then $\psi=c+v$ for some $c\in S^{1}$, $v\in\dot{H}^{1}({\mathbf{R}}^{3})$. For $d=2$, unbounded phase oscillations may occur at spatial infinity that rule out characterizing the energy space with an affine structure. The space $(\mathbb{E}({\mathbf{R}}^{2}),d_{\mathbb{E}})$ is not separable. Given its relevance for the well-posedness theory, this question is going to be addressed in detail in Section 2. In particular, one may introduce a weaker topology that restores separability and connectedness. Note that this affine structure of the energy space is available for higher dimensions $d\geq 4$ to which our approach adapts. As $\mathbb{E}({\mathbf{R}})\subset X^{1}({\mathbf{R}})$, the local well-posedness theory simplifies for $d=1$. We expect our approach to extend to $d=1$. Previous results [24, 25, 28] do not cover the full generality of Assumption 1.1.
@@
Namely, for any $r,R>0$ and $\psi_{0}^{\ast}\in\mathbb{E}({\mathbf{R}}^{d})$ such that $\mathcal{E}(\psi_{0}^{\ast})\leq R$ let $\mathcal{O}_{r}:=\{\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d}):d(\psi_{0},\psi_{0}^{\ast})\leq r\}$. Then, there exists $T^{\ast}(\mathcal{O}_{r})>0$ such that $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{d}))$ for all initial data $\psi(0)=\psi_{0}\in\mathcal{O}_{r}$. Moreover, for any $0<T<T^{\ast}(\mathcal{O}_{r})$ there exists $C>0$ such that for any $\psi_{1},\psi_{2}\in C([0,T];\mathbb{E}({\mathbf{R}}^{d}))$ with initial data $\psi_{0}^{1},\psi_{0}^{2}\in\mathcal{O}_{r}$, we have
@@ 1.3. Global well-posedness results
This assumption yields that $F$ achieves a local minimum for the constant solution $|\psi|^{2}=1$. In nonlinear optics, his requirement appears in the physical literature to prevent the onset of modulational instability, also known as Benjamin-Feir instability [6] of the constant equilibrium solution, i.e. the continuous wave background [48, 62].
@@
with $a>0$, $0<\alpha_{1}<2$ and where $g$ satisfies Assumptions 1.1 (K1) and (K2) for some $0\leq\alpha_{2}<\alpha_{1}$. In addition, $F$ is such that $F(\rho)>0$ for all $\rho>1$.
@@
The assumption on the roots of $F$ allows for physically relevant nonlinearities to be studied. It appears from the physics literature [5, 49, 62] that in relevant applications the largest root of $F$ corresponds to the far-field behavior $\rho_{0}=1$ and constitutes a local minimum of $F$ which is linked to preventing modulational instability of the continuous background wave [48, 62]. To obtain global existence, we rely on the aforementioned affine structure of the energy space $\mathbb{E}({\mathbf{R}}^{3})$ and require that $\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{3})$ while $v_{0}\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ only yields $|v_{0}|^{2}+2\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{3})$. An exponential bound on $\operatorname{Re}(\overline{c}v)(t)\in L^{2}({\mathbf{R}}^{3})$ is derived which compensates for the lack of the conserved mass due to the non-trivial farfield. The result of Theorem 1.7 remains valid if the assumption $\operatorname{Re}(\overline{c}v_{0})\in L^{2}$ is replaced by a smallness assumption on $\mathcal{H}(\psi_{0})$ and $\|\nabla\operatorname{Re}(\overline{c}v_{0})\|_{L^{2}}^{2}$ depending only on the second largest positive root of $F$, see Remark 4.12.
Finally, we consider the general scenario in which $F$ satisfies Assumption 1.5 is satisfied but may be unbounded from below, e.g. in the case of competing power-law nonlinearities with the focusing one dominant at large intensities.
@@
To the best of the authors’ knowledge, global results for (1.1) in $d=2,3$ are in general not available in the literature in the case of non sign-definite total energies unless the nonlinear potential energy density is assumed to be bounded from below and in addition more regularity on $f$ [25], or in the cubic-quintic case a condition on $\operatorname{Re}(v)$ like the ones mentioned are assumed, cf. [47, 45] are assumed. In [46] small-data global well-posedness for cubic-quintic nonlinearities is proven also in the case where the quintic nonlinearity is focusing. In [59] the authors consider for $d=1,2$ nonlinear potentials energies unbounded from below for specific regular energy subcritical nonlinearities and prove small data global existence for solutions of the form $\psi=1+u$ in tailored function spaces.
The main steps of our approach are briefly sketched. First, we identify the suitable mathematical setting for our analysis, namely the energy space $\mathbb{E}$, see (1.9). We crucially rely on the fact that $(\mathbb{E},d_{\mathbb{E}})$ is a complete metric space as well as the properties of the free propagator introduced in [26, 27]. The Hamiltonian $\mathcal{H}$ is well-defined for functions in $\mathbb{E}$. While wave-functions in $d=3$ can be decomposed as $\psi=c+v$ with $|c|=1,c\in{\mathbf{C}}$ and $v\in\dot{H}^{1}({\mathbf{R}}^{3})$, for $d=2$ the wave-functions may exhibit unbounded oscillations of the phase at spatial infinity. This motivates treating separately the well-posedness problem for $d=2,3$. In both cases, we show local existence of a solution in the affine space $\psi=\psi_{0}+H^{1}({\mathbf{R}}^{d})$ by a perturbative Kato-type argument [42] and also [14, Chapter 4]. Subsequently, uniqueness in $C([0,T];\mathbb{E}({\mathbf{R}}^{d}))$ is proven. The fixed-point argument only provides continuous dependence with respect to perturbations in the space $\psi_{0}+H^{1}({\mathbf{R}}^{d})$. The proof of continuous dependence on the initial data with respect to the topology induced by the metric $d_{\mathbb{E}}$ requires additional estimates and differs in a substantial way from the $H^{1}$-well-posedness theory for NLS-equations with vanishing conditions at infinity. This is due to the non-integrability of wave-functions and the intricate topological structure of the energy space linked to the far-field behavior including oscillations of the phase and the low regularity of the nonlinearity. Global well-posedness is shown relying on the conservation of the Hamiltonian $\mathcal{H}$.
@@
For the 3D energy-critical quintic equation, one may proceed as described in Section 1.4.
@@
Beyond the mentioned power-type nonlinearities, here are some examples of physically relevant nonlinearties and far-field (1.2):
@@
    saturated nonlinearities $f(\rho)=\frac{\rho}{1+\gamma\rho}-\frac{1}{1+\gamma}$ with $\gamma>0$, see for instance [65, Chapter 9.3] and references therein,
@@
    exponential nonlinearities $f(\rho)=(\mathrm{e}^{-\gamma}-\mathrm{e}^{-\gamma\rho})$ with $\gamma>0$ [65, Chapter 9.3],
@@
    transiting nonlinearities of the form $f(\rho)=2\rho\left(1+\alpha\tanh\left(\gamma(\rho^{2}-1\right)\right)$ occurring in nonlinear optics [62, Section VI],
@@
The cubic-quintic equation (1.18) falls within (1) of the aforementioned list and is also recovered in the small amplitude approximation of (2) and (3) of the above examples [65, Chapter 9.3]. We also refer to [18] for a more detailed overview.
Remark 1.10.
Nonlinear Schrödinger equations equipped with non-vanishing boundary conditions (1.2) are also investigated with non-local interaction, e.g. in the context of supersolids, see [55] for an overview of relevant models. While we do not pursue this topic here, we expect that our analysis combined with [23] allows for a straightforward extension to general non-local nonlinearities.